A company acquired a new piece of equipment on January 1, 2011 at a cost of $200,000. The equipment is expected to have a useful life of 10 years, a residual value of $20,000 and is depreciated on a straight-line basis. On January 1, 2013, the equipment was appraised and determined to have a fair value of $190,000 and a residual value of $25,000 and a remaining useful life of 10 years.

At what amount should the equipment be reported on the December 31, 2013 balance sheet under the IFRS revaluation model?

$136,000

$173,500

$110,000

$190,000

$165,000
Answer #1

Jan 1 valuation = 190,000

Depreciation for 2013 = (190000-25000)/10 = 16500

Value on December 31 = 190000-16500 = 173500

Answer is:

173500

Please rate